UT November 28 2018

LBTO Engineering Night

OSA:. D.G.Huerta

ISp: E. Solheid

ISA: B. Rothberg (Tucson)

AO Support: G. Taylor

Software Support: P. Kubanek

Observers: B. Rothberg, O. Kuhn, J. Hill

Summary:

1) Collected MODS-Bino transform data for D. Miller (who will process it ASAP) - however new results are not conclusive and a further updated transform won't be available until next day.

Tests of IT 7387, 7218

2) Tested monocular LBC-R guiding - it works ok.

3) Tested mixed-mode (MODS+LBC-R) in binocular and pseudo-monocular. LBC-R fails in binocular mode after first dither position with "CS IIF_GetRotatorTrajectory error: Timed out waiting for target OK" and "QueryRotator failed, trajectory is invalid", LBC-R in pseudo-monocular now does NOT work with any dithering, even when first dither position is 0, 0.

Conclusion - pseudo-mono MODS-LBC no longer works, Bincoular MODS-LBC - no longer works

4) NS guiding testing with LBC-R monocular. Works with high-speed (-160" /hr RA 42"/hr Dec) for 60sec and 180sec integrations.

5) Given the variable seeing conditions, the oncoming snowstorm this weekend, and that we had exhausted the transform and mixed mode tests, we decided to switch to backup D-time programs - two targets observed successfully.

LUCIs were not used tonight.

Pointing with MODS seems better than with LUCI over the last week


UT 12:48 - LBCs turned on, MODS awake, LUCI software up

UT 12:50 - Init all on LUCIs run, two test darks taken each LUCI

L1 0001-0002; L2 0001-0002

- had a number of alarms go off - HBS went off, high pitched alarm went off, DG. Huerta working on recovery - possible brownout?

UT 01:17 - collimating on-sky

UT 01:34 -D.G. Huerta and G. Taylor working on AO Checkout - having some issues.

UT 01:47 - mountain comm drop out

UT 01:58 - still working on AO checkout

UT 02:02 - AO checkout complete, reconfiguring for binocuar MODS, bounce PCS

UT 02:12 - going to do Stone A1 at PA=180

UT 02:19 - setting up MODS for transform collection - transforminit_PointingMask.pro - mods script located in ~modeseng/modsScripts/Support

sets up 1kx1k, puts in pointing mask, 10sec each side (for now)

UT 02:30 - pointing and collimation check

UT 02:49 - having trouble collimating - clodus rolled in fast

UT 02:59 - failed to find the guidestar for Stone A1 - looking for brighter stars (initial collimation was R=10th mag so we are working with thick clouds here).

UT 03:05 - Olga revised the list for brighter stars - giving it a shot

upping exposure time to 30seconds, manually running IMCS lock first

mods1r.20181128.0002 - test exposure of star

mods2b.20181128.0002 - test exposure of star

Can't see a 10th mag star in 30seconds with MODS.....so.......yeah....clouds

1R 0003, 2B 0003 - see a star - very faint (peak ~700 counts)

UT 03:28 - new brighter saturated .... thnk clouds are clearing so going back to original Stone A1 star list for transform

UT 03:32 - drop out communications to the mountain

- resending preset because the guideprobe was commanded to move before guiding was paused and this all occurred same time as com dropout

- looks much better - clouds may have rolled out for a while now

UT 03:38 - drop out communications to the mountain

- Paused Guiding

1R 0005 - IMCS lock was not on - double image

2B 0005 - IMCS lock was not on - double image

1R 0006 - 30sec

2B 0006 - 30sec

1R 0007 - 60 sec - looks good

2B 0007 - 60 sec - looks good

UT 03:49 - starting transform collect

1R 0008; 2B 0008 - start of trasnform data set

- Guiding was paused when we started

-Left seems to have recovered, need to restart on right hand side

UT 03:52 - resending preset on right side

UT 03:52 - we need to check pointing - so stopping (no data taken)

UT 03:58 - pointing check complete, stopping and starting MODS AGW services

Stone A1

UT 03:59 - sending preset

UT 04:02 - starting transform collection

first exposures 1R 0008 ; 2B 0008

0008,0009 - don't see a star, don't see all of the circles - probe must have been blocking the 30" FOV

0010 - can finally see stars

0011 - background image

UT 04:16 - MODS-2 failed on first star, MODS-1 failed on first star

restarting with star #2

canceled preset

IDL is acting odd, it seems to want to send preset rather than set them up, exiting both IDL sessions, starting over

UT 04:25 - resending preset

/home/lbto/data/20181128/Stone_A1_star15_pa180.list

0010 - On Axis

0011 - Background

0012 - GS1

0013 - GS2

0014 - GS3

0015 - GS4

0016 - GS7

0017 - GS8

0018 - GS9

-last guidestar - barely collimating on it

0019 - GS16

0020 - On Axis

0021 - BIAS

UT 05:03 - completed transform

MODS Transform Analysis -Doug

MODS1

Determine direction of motion of target on MODS1 due to an guide probe motion

TCS log of a offset in MODS1 imaging mode
2018-11-27T08:19:05.610081+00:00 tcs1 LBT_GCSL: AGW_MODS: doSetProbePosition.  New position: -34.239, -142.085 OnAxisPosition: 0.000, 0.000
2018-11-27T08:19:17.293851+00:00 tcs1 LBT_GCSL: AGW_MODS: doUpdateProbePosition: mods_getxy(192.168.139.130,0,..) returned 34.7040, -142.0920 AGw coords
2018-11-27T08:24:33.725914+00:00 tcs1 LBT_GCSL: AGW_MODS: doSetProbePosition.  New position: -34.525, -135.550 OnAxisPosition: 0.000, 0.000
2018-11-27T08:24:39.152427+00:00 tcs1 LBT_GCSL: AGW_MODS: doUpdateProbePosition: mods_getxy(192.168.139.130,0,..) returned 34.9690, -135.5610 AGw coords

Focal Plane (PCS Coordinates)
  • X: -34.525 - (-34.239) = -0.286003 mm
  • Y: -135.550 - (-142.085) = 6.53500 mm

Guide Probe
  • X: 34.9690 - 34.7040 = 0.265003 mm
  • Y: -135.5610 - (-142.0920) = 6.53099 mm

Image Motion (mods1r.20181127.0010.fits - mods1r.20181127.0009.fits collected on 20181127)
  • X: 501 - 505 = -4 pix
  • Y: 621 - 531 = 90 pix

Relationship between Image Motion and Guide Probe Position
  • Checked that the same target and guide star coordinate were used in both images
  • X: negative (-) image motion = Positive (+) guide probe motion (opposite sign => transform_params[i].x_flip = 1)
  • Y: positive (+) image motion = positive (+) guide probe motion ( same sign => transform_params[i].y_flip = 0)

  • MODS1 Image -vs- Probe Motion:
    mods1r_image-v-probe_motion.jpg

MODS2

Determine direction of motion of target on MODS2 due to an guide probe motion

2018-11-28T13:07:50.244399+00:00 tcs1 LBT_GCSR: AGW_MODS: doSetProbePosition.  New position: -10.814, -104.675 OnAxisPosition: 0.000, 0.000
2018-11-28T13:07:56.618846+00:00 tcs1 LBT_GCSR: AGW_MODS: doUpdateProbePosition: mods_getxy(192.168.139.230,1,..) returned 9.4570, -104.7930 AGw coords
2018-11-28T13:13:08.794102+00:00 tcs1 LBT_GCSR: AGW_MODS: doSetProbePosition.  New position: -8.571, -99.099 OnAxisPosition: 0.000, 0.000
2018-11-28T13:13:14.228232+00:00 tcs1 LBT_GCSR: AGW_MODS: doUpdateProbePosition: mods_getxy(192.168.139.230,1,..) returned 7.2900, -99.2300 AGw coords

Focal Plane (PCS Coordinates)
  • X: -8.571 - (-10.814) = 2.24300
  • Y: -99.099 - (-104.675) = 5.57600

Guide Probe
  • X: 7.2900 - 9.4570 = -2.16700
  • Y: -99.2300 - (-104.7930) = 5.56300

Image Motion (mods2b.20181120.0035.fits - mods2b.20181120.0034.fits collected on 20181127)
  • X: 510 - 482 = 28
  • Y: 590 - 512 = 78

Relationship between Image Motion and Guide Probe Position
  • Checked that the same target and guide star coordinate were used in both images
  • X: positive (+) image motion = negative (-) guide probe motion (opposite sign => transform_params[i].x_flip = 1)
  • Y: positive (+) image motion = positive (+) guide probe motion ( same sign => transform_params[i].y_flip = 0)

  • MODS2 Image -vs- Probe Motion:
    mods2b_image-v-probe_motion.jpg


UT 05:04 - Reconfiguring for MODS-1 SX, LBC-R DX

Putting MODS2 to sleep for NOW

First test will be LBC-R only guiding - IC 342 - 5pt dither, 50", 4 filters

UT 05:35 - sending CP IC342 script - will dohybrid

UT 05:52 - colliation done, moving to IC 342

UT 06:27 - script for IC 342 done

UT 06:28 - moving to NGC 1068 collimation field

UT 06:31 - running dofpia

UT 06:36 - authorizing MODS-1 SX + LBC-R binocular

-checking MODS-1 pointing and collimation

UT 06:44 - sending binocular LBC-R and MODS-1 presets

....as a scoundrel once said ..."Watch this..."

UT 06:47 - LBC-R is taking exposures.... first dither is -5 10 DX DY, MODS-1 is happily proceeding with acquisition

UT 06:53 - acquisition complete - spectra started

UT 06:54 - "Get Trajectory Failed" LBC - timed out waiting for target

-this looks like an LBC-Red rotator failure unrelated to the binocular test

UT 06:58 - Stopping MODS and LBC-R - resending both presets

UT 06:59 - LBC-R started exposure

Where's the hydrospanner....

UT 07:02 - happened again - 07:02:25 QueryRotator failed, trajectory is invalid -

before that TCS IIF_GetRotatorTrajectory error: Timed out waiting for target OK

LBC-R completes the first two exposures at dither position one, then attempts to move and fails

Going to re-authorize for psuedo-monocular - MODS in the driver seat

UT 07:09 - running PM - aligning MODS-1 first

UT 07:17 - starting MODS spectra, starting LBC - first dither in LBC script is -5,-10

LBCR.20181128.071752 shows a double image (F972N20 - 1st exposure) - taken during alignment

071929 - 2nd exposure in series OK (Y-Fan)

072057 - 1st exposure in offset (F972N20) jumped ~50 pixels (~11") - MODS-1 was stationary - both LBC jumps are vertical (DET Y)

072400 - no jump F972N20 (3rd dither, 1st exposure)

072713 - jump (F972N20) - 4th dither, 1st exposure

073025 - jump (F972N20) - 5th dither, 1st exposure

-always seems to be first LBC exposure after dither

UT 07:40 - created a 3fiilter version of LBC-R pseudo script, z-sloan is now first filter, will run this and MODS stare script again, 3filter LBC script has 0,0 as first dither position

UT 07:41 - started new test

UT 07:43 - DX exceeding forces

074159 - z-sloan first in dither sequence jumped - confirmed on LBTplot of LBC guiding

074338 - F927N20 - 2nd in dither sequence jumped - but this could be a guiding issue (guidestar too faint?) the jump is NE not vertical

074515 - Y-Fan also jumped but elongated diagonally

074633 - also jumped - but multiple times - errors in LBC logs says no good source found - this may be the source of the non-vertical jumps

AllowOffset radio button is checked in IIFGUI on right side - so this is NOT the issue/cause

UT 07:53 - stopping LBC-R script, DX active optics being cleared -

UT 07:55 - stopped MODS script too

M1 DX is still reaching limits - pointing offsets were only 5-10" on DX

UT 07:56 - moving to zenith to fix issues with unhappy DX primary

-NOTE: First dither for LBC in pseudo-mono mode WAS 0,0

UT 08:10 - we are going to test NS guiding using C. Veillet's 2nd target.....

UT 08:14 - authorizing LBC-Red only

UT 08:18 - sending collimation to filed near NS target

UT 08:21 - running dofpia

TARGET is K18W00R

UT 08:27 - collimated

UT 08:28 - sending preset - 5pt dither -

-see guidestar in box in 1st and 2nd position

every star is streaked 17 pixels with an upturn

streak should be ~2.8" in RA and 0.8" i Dec (each exposure is 60sec)

UT 08:38 - testing version with 180sec integration time -

083915 - guiding failed after 85sec ... suppose the guidestar flew out of the field

2nd image guiding failed after ~50sec

084634 - 3rd image failed at 90sec guiding - looks like a star at the edge of the guider (see it in cutout in GUI) - multiple jumps - not as good as previous images

UR 08:58 - completed NS tests

Testing Mono-LBC-R guiding works ok - Perseus A

UT 09:08 - collimating near Perseus A

DIMM 1.5"

UT 09:10 - moving to Perseus A

UT 09:19 - data looks good

UT 09:38 - script done

UT 09:41 - reconfiguring for MODS-Binocular

Reverting PCSInstrument.conf back to zp RDG = 27.55 and reverting AGW.ini on MODS2 computer since new transform not ready yet

UT 10:06 - pointing and collimation check

-preset hung and SX GCS hung - reauthorized

UT 10:12 - restarting GCS left - worked

UT 10:28 - sending bino preset to SDSSJ0849+1114

-need to move MODS2B over by 2 pix

UT 10:44 - starting spectroscopic exposures

seeing seems variable - hope for the best

UT 10:54 - WFS on both showing 0.9-1.2"

UT 11:09 - taking another 1200sec exposure - lots of glints in Blue exposures (close to Moon)

-second exposure the seeing on DIMM and WFS has been < 1"

UT 11:30 - taking 3rd exposure

UT 11:38 - mountain com drop out

mods1r 0036-0041

mods1b 0008-0010

mods2b 0022-0028

UT 11:53 - Moving to MS0754.6+3928 - QSO

UT 12:08 - starting spectroscopy

DIMM is 1.5" WFS 1.2-1.5" - this target is a poor conditions backup target

Taking a 2nd 1200sec exposure

UT 13:04 - can see thin wisps of clouds overhead as illuminated by Moon

mods1r 0042-0046

mods1b 0011-0012

mods2b 0029-0033

UT 13:07 - moving to Feige 34 - specphotstd

UT 13:15 - starting spectroscopic exposures

mods1r 0047-0051

mods1b 0013-0015

mods2b 0034-0038

UT 13:29 - closing up - will take MODS-2 cals, MODS-1 cals will be used from last night, no LUCI cals, will take some LBC-R biases

UT 13:40 - starting MODS2 cals, LBC-R biases (10)

UT 13:42 - putting MODS-1 to sleep

UT 13:48 - powered off LBC-R (LBC-Blue never came up)

mods2b 0039-0041 - Arc Lamps

0042-0051- -slitless pixel flats

0052-0056 - 8k Biases

UT 14:22 - putting MODS-2 to sleep

ALL DONE!

-- BarryRothberg - 28 Nov 2018
Topic attachments
I Attachment Action Size Date Who Comment
mods1r_image-v-probe_motion.jpgjpg mods1r_image-v-probe_motion.jpg manage 254 K 03 Dec 2018 - 18:06 DougMiller MODS1 Image -vs- Probe Motion
mods2b_image-v-probe_motion.jpgjpg mods2b_image-v-probe_motion.jpg manage 250 K 03 Dec 2018 - 19:10 DougMiller MODS2 Image -vs- Probe Motion
Topic revision: r4 - 03 Dec 2018, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback